OpenStudy (anonymous):

the product of the square of a number and 5

OpenStudy (dangerousjesse):

\(\sqrt{x\times 5}\) or \(\sqrt{x}\times 5\)

OpenStudy (anonymous):

that would be the square root!

OpenStudy (anonymous):

if you call the number \(x\) then its square is \(x^2\)

OpenStudy (dangerousjesse):

I misread, sorry. \(x^2\times 5\)

OpenStudy (anonymous):

hi brother :)

OpenStudy (anonymous):

@nirmalnema

OpenStudy (nirmalnema):

hiiii...

OpenStudy (anonymous):

sqrt(n) - (5n) [sqrt(n)]^2 - (5n)^2 n - 5^2 * n^2 Answer: n - 25n^2

OpenStudy (anonymous):

hw r u brother?

OpenStudy (nirmalnema):

im gud aap batoo

OpenStudy (anonymous):

i am also good :)

OpenStudy (nirmalnema):

ok.. :)